### v2

@@ Results
After the models are fit on the datasets, their performance is evaluated using the accuracy metrics from Section 4 and their values are reported in Table 2. Along every metric (column), each model is colored by rank with the best model colored purple and the worst model colored red. Cells colored black correspond to values that diverged and are automatically assigned last place. These individual ranks are then summed to produce a final model score which is used to sort the table and quantify relative performance.
@@
Table 2 should be read as follows: The first column corresponds to the specific model that was regressed, and the subsequent two columns define the data conditions used for the regression. The $N$ column corresponds to the number of training data and the error column corresponds to the magnitude of the random of the acceleration vector (i.e. either 0% or 10%). When these cells are colored light red, that means the condition is the less desirable configuration and is used to stress model performance, whereas white corresponds to the favorable data condition.
@@
The fourth column is the aforementioned model score which sums the metric ranks for the listed model. For example, the small PINN-III (PINN-III S) is ranked 4th on the planes metric, 5th on the interior, 5th on exterior, 3rd on extrapolation, 9th on surface, and 3rd on the trajectory metric. Added together, these sum to a model score of 29 which is the best score among the 62 tested models. The individual ranks for every model in each metric are reported separately in Appendix D Table 5 for reference.
@@
Table 2 shows that the three highest scoring gravity models are the PINN-III S, PINN-III L, and Mascon L models respectively. Each of these models are regressed under the favorable 50,000 and 0% error data conditions, and they all maintain low error across the different metrics. The PINNs generally perform better near the surface of the asteroid, where the mascon model is known to struggle; however, the mascon model performs better in the high-altitude extrapolation regime where no training data was present.
@@
Notably, six of the eight PINN-III configurations score in the top ten models, despite some of these models being trained under noisy or sparse data conditions. For example, the 4th best model is a PINN-III S, but this model is only trained with 500 data points. The next best model trained under these conditions is a Mascon L model which scores 8th and uses a two orders of magnitude more parameters. Similarly a PINN-III L performs the best among all models trained with 10% error on the acceleration vectors — scoring 7th across all models. Finally, a PINN-III L is also the best performer of all models trained on both sparse and noisy data, scoring in the upper 35% percentile and placing 22nd total. Table 3 sorts and ranks these models by their respective training conditions for a more compact summary of model performance and again highlights how PINN-III S and L remain the two highest scoring models across all tested conditions.
@@
Table 2 simultaneously illustrates the accuracy and robustness of the PINN-GM-III to these stressful data conditions while also exposing the brittleness of other gravity models — including past PINN-GM generations. For example, PINN-II L has the lowest error at the surface of any gravity model, averaging 2.8% across all 200,700 facets on the shape model. However, when tested in the extrapolation regime, PINN-II diverges and ranks last of all 62 models. Similarly, the Mascon L has the highest score of all analytic models when regressed on ideal training conditions, but when trained on sparse and noisy data sets, this model does not remain robust — ranking 60th of 62.
Past machine learning models have even more difficulty, with the traditional neural networks, ELM, and PINN-I consistently scoring in the lower 50th percentile. These struggles highlight just how data intensive the original machine learning implementations can be. Even with 50,000 noiseless data points, these model are unable to converge — illustrating the need for more nuanced discussion about how and when these models can break down. While the regression strategies used to fit these models sought to remain close to the original implementations, it remains possible that different training configurations and practices could yield better performance.
@@
PINN III L 1 3 1 2
PINN III S 3 1 2 1
PINN II S 2 5 4 4
PINN II L 4 4 5 5
MASCONS L 15 2 3 3
@@
MASCONS S 8 6 6 8
PINN I L 5 9 7 9
POLY L inf inf inf 7
SH S 6 7 9 11
PINN I S 7 8 8 10
PM - 9 10 10 12
SH L 10 11 11 13
ELM L 12 13 14 16
TNN S 13 14 13 14
ELM S 12 13 15 17
TNN L 14 15 12 15
Table 3: Gravity models ranked and scored via evaluation metrics
@@
While Tables 2 and 3 highlight the relative accuracy of the available gravity models, another important comparison point is each model’s evaluation speed. To characterize this, 1,000 randomly distributed test points are evaluated using each gravity model. The total evaluation time is measured and an average inference time per sample is calculated and reported in Figure 7.
@@
For the large models, Figure 7 shows that the polyhedral model of 30,006 parameters — which corresponds to approximately 10,000 facets — is the slowest gravity model averaging over 20 ms for each evaluation. The PINN-GM-III and II are an order of magnitude faster at approximately 2 ms, followed by the mascon model at 0.8 ms. For small models, the PINN-III is the slowest at 0.9 ms, followed by the remaining machine learning gravity models. The small polyhedral model (204 parameters or 66 faces) is much faster at 0.1 ms. Finally, the spherical harmonic and point mass are the fastest, executing in less than 0.01 ms per evaluation.
These results highlight that speed is among the largest trade-offs for the PINN-GM-III, particularly in the small model regime. That said, the PINN-III models are not prohibitively slow, and their computational cost remains somewhat constant irrespective of model size. For example, PINN-III L is two orders of magnitude larger than PINN-III S, but its computational cost is only twice as large. In contrast, the large polyhedral models cost appears to scale linearly, taking approximately 100 times longer for a model that is 100 times bigger. This suggests that the PINN-III remains a strong option if robustness and accuracy are the primary goal, and users can increase the size and performance of these models with relatively little added overhead.
@@ 6 Conclusions
Scientific machine learning and physics informed neural networks offer a compelling set of tools to address the gravity field modeling problem. Rather than using prescriptive analytic gravity models prone to various limitations, PINNs can learn convenient representations of the gravitational potential while maintaining desirable physics properties and assurances. While past machine learning gravity models have offered early glimpses into the potential advantages of these numerical solutions, this greater class of model remained prone to multiple pitfalls that had yet to be exposed or addressed. This paper highlights these failure cases, and introduces design modifications within the machine learning architecture that can overcome these challenges. Taken together, these modifications form the third generation PINN gravity model, or PINN-GM-III. This model is designed to solve the problems of feature divergence, bias towards low-altitude samples, numerical instability, and extrapolation error, while also proposing a framework for fusing analytic and numerical gravity models together for enhanced modeling accuracy. While these modifications are studied exclusively on the PINN gravity model, it should be noted that many of these modifications can be applied to other machine learning solutions to enhance their own robustness and generalizability.
Beyond introducing the PINN-GM-III, this manuscript also proposes new evaluation metrics to more comprehensively assess the modeling accuracy of various analytic and numerical gravity models in different orbital regimes. When trained on data from a heterogeneous density asteroid, the PINN-GM-III is shown to achieve competitive performance over past generations and other numerical and analytic models, while also demonstrating robustness to data sparse and noisy conditions. Future work will continue to investigate design modifications that can improve model performance, with a particular emphasis on returning to large celestial bodies and investigating ways in which the high-frequency components can be learned and represented more efficiently.
@@ Appendix A PINN-GM-III Training Details
Appendix B Ablation Studies
@@ Appendix C Comments on Past Machine Learning Performance
For GeodesyNets izzoGeodesyIrregularSmall2022 , SIRENs of 9 hidden layers with 100 nodes each are used ($\approx 100^{2}*(9-1)+8*100=80,800$). Four asteroids are studied: Bennu, Churyumov-Gerasimenko, Eros, and Itokawa. In their supplementary materials (Table S4), the relative error about Eros is reported at three characteristic altitudes. At their lowest altitude, the average error is 0.571% and their highest altitude is 0.146%. In attempts to quantify error across the entire high and low altitude regime, these values are averaged for the reported value of 0.359%.
@@ D.5 Neural Network Models
D.6 Comparison Study Model Ranks
Table 5 reported the exact value for each metric for every regressed model and their rank was expressed via color in the table. Here the exact rank is reported for greater transparency.

### v3

@@ Results
After the models are fit on the datasets, their performance is evaluated using the accuracy metrics from Section 4 and their values are reported in Table 2. Along every metric (column), each model is colored by rank with the best model colored purple and the worst model colored red. Cells colored red or black correspond to values that diverged and are automatically assigned last place (70th). In the case of the planes, interior, exterior, extrapolation, and surface metrics, this divergence corresponds with values that exceed 100% error and therefore are not usable. These individual ranks are then summed to produce a final model score which is used to sort the table and quantify relative performance.
@@
Table 2 should be read as follows: The first column corresponds to the specific model that was regressed, and the subsequent two columns define the data conditions used for the regression. The $N$ column is the number of training data and the error column is the magnitude of the acceleration vector error (i.e. either 0% or 10%). When these cells are colored light red, that means the condition is less desirable and is used to stress model performance, whereas white corresponds to the favorable data condition.
@@
The fourth column is the aforementioned model score which sums the metric ranks for the listed model. For example, the small PINN-III (PINN-III S) is ranked 4th on the planes metric, 5th on the interior, 4th on exterior, 3rd on extrapolation, 8th on surface, and 5rd on the trajectory metric. Added together, these sum to a model score of 29 which is the second best score among the 70 tested models. The individual ranks for every model in each metric are reported separately in Appendix D Table 6 for reference.
@@
Table 2 shows that the three highest scoring gravity models are the Mascon L, PINN-III S, and PINN-III L models respectively. Each of these models are regressed under the favorable 50,000 and 0% error data conditions, and they all maintain low error across the different metrics. The PINN-III L performs better near the surface of the asteroid, where the mascon model is known to struggle; however, the mascon model performs better in the high-altitude extrapolation regime where no training data was present.
@@
Notably, five of the eight PINN-IIIs score in the top ten models, despite some of these models being trained under noisy or sparse data conditions. For example, the 5th best model is a PINN-III S, but this model is only trained with 500 data points. The next best model trained under these conditions is a Mascon L model which scores 9th and uses a two orders of magnitude more parameters. Similarly a PINN-III L performs the best among all models trained with 10% error on the acceleration vectors — scoring 6th across all models. Finally, a PINN-III L is also the best performer of all models trained on both sparse and noisy data, scoring in the upper 30th percentile and placing 18th overall. Table 3(a) sorts and ranks these models by their respective training conditions for a more compact summary of model performance and again highlights how PINN-III S and L remain the two highest scoring models across all tested conditions on a heterogenous density Eros.
@@
Table 2 simultaneously illustrates the accuracy and robustness of the PINN-GM-III to these stressful data conditions while also exposing the brittleness of other gravity models — including past PINN-GM generations. For example, PINN-II L has the lowest error at the surface of any gravity model, averaging 2.5% across all 200,700 facets on the shape model. However, when tested in the extrapolation regime, PINN-II diverges. Similarly, the Mascon L has the highest score across all models when regressed on ideal training conditions, but when trained on sparse and noisy data sets, this model diverges in all but one metric, ranking 58th of 70 overall.
Past machine learning models exhibit greater concern, with the traditional neural networks, ELMs, PINN-I, and GeodesyNets consistently scoring in the lower 50th percentile. The candidate failure modes discussed in Section 3 are best captured by inspecting the interior, exterior, and extrapolation metrics. Note how these models exhibit deteriorating accuracy when evaluated on increasingly high altitude data — i.e. the interior values are generally lower than exterior values. Moreover, these vast majority of models — GeodesyNets excluded — meet the divergence criteria when tested beyond the bounds of the training data. These behaviors provide confirmation of the challenges many machine models face when tested across a more diverse set of experiments, and the PINN-GM-III avoids these same challenges due to the new suite of proposed modifications.
As a caveat, GeodesyNets are not prone to the exact same failure modes as the other machine models due to their fundamentally different architecture and evaluation procedure. Because the densities predicted are only computed within a unit volume, these models do not pose the same risk of feature or model divergence at high altitudes. That said, these models still struggle to converge as there does not exist sufficient amounts of data to reliably constraint the density function. Even with 50,000 noiseless data points, the GeodesyNets yield errors in excess of 80% across the different testing regimes. These results further highlight how sensitive and data intensive these alternative ML model implementations can be — illustrating the need for more nuanced discussion about how and when these models can break down. While the regression strategies used to fit these models sought to remain close to the original implementations — as is discussed in Appendix D — it remains possible that different training configurations and practices could yield better performance.
To ensure the conclusions drawn through this study are generalizable, three additional gravity fields are tested. These analyses regress the same 70 models on a homogeneous density version of Eros, and a homo- and a heterogenous density version of the asteroid Bennu. The imposed heterogenous density profiles again remain ±$\pm 10\%$ mass at locations $\pm$±R$/2$ along the x-axis of the respective asteroid. These four environments attempt to span common and exotic gravity fields: from nearly spherical to irregular geometries, and from homogeneous to heterogeneous densities profiles. By studying the performance across these different cases, the general behaviors of these models can be reasonably inferred for other candidate small bodies. The exact metric values for these additional cases are reported in Appendix D.7, and their score summaries across the different data conditions are reported in Table 3.
Across these three additional experiments, the PINN-GM-III remains competitive — consistently reporting the highest score among the machine learning models. Notably, the new scenarios do show the strength of the analytic models, such as the spherical harmonic model for the asteroid Bennu, and the polyhedral model for the constant density cases. Importantly, the results for the polyhedral model are not directly comparable to the other models reported, as the polyhedral shape was not constructed / regressed from the position and acceleration data. Its inclusion can only be used to infer the strength of the model in a parametric sense, where the polyhedral models of approximately 30,000 parameters achieves better performance than other models of the same size. Regardless, these results imply that the PINN-GM-III exhibits its greatest utility when modeling asteroids with irregular shapes and unknown density.
@@
PINN III L 1 2 1 3
PINN III S 2 1 6 2
PINN II L 3 4 3 4
MASCONS L 15 3 2 1
PINN II S 4 6 5 5
MASCONS S 9 5 4 7
@@
PINN I L 6 13 7 9
PM - 5 7 10 12
POLY L inf inf inf 8
SH S 7 8 9 10
PINN I S 8 9 8 11
GEONET L 10 10 12 14
SH L 11 12 11 13
GEONET S 12 11 13 15
TNN S 17 16 14 16
ELM L 13 14 16 18
ELM S 14 15 17 19
TNN L 16 17 15 17
Eros Heterogeneous
N 500 50000
Error (%) 10 0 10 0
Model
POLY L inf inf inf 1
PINN III S 2 2 1 4
PINN III L 2 2 2 3
PINN II L 3 4 5 6
POLY S inf inf inf 5
PINN II S 4 6 6 7
MASCONS L 15 3 3 2
MASCONS S 8 5 4 8
PM - 5 7 9 12
PINN I L 7 13 7 9
PINN I S 9 11 8 10
SH S 7 8 10 11
GEONET L 10 9 11 13
GEONET S 12 11 13 15
SH L 11 12 12 14
TNN S 16 16 14 17
ELM L 13 14 16 18
ELM S 14 15 17 19
TNN L 17 17 15 16
Eros Homogeneous
N 500 50000
Error (%) 10 0 10 0
Model
SH S 1 1 1 4
PINN III L 2 2 4 2
PINN III S 5 3 7 3
PINN II L 3 4 6 5
PINN II S 6 6 3 6
MASCONS L 15 5 5 1
PM - 4 7 9 11
MASCONS S 14 10 2 7
PINN I L 7 8 8 10
POLY L inf inf inf 8
PINN I S 9 12 10 12
POLY S inf inf inf 10
SH L 8 9 11 13
GEONET L 10 11 12 14
GEONET S 11 13 13 15
TNN L 16 17 14 16
ELM L 12 15 16 18
ELM S 14 14 17 19
TNN S 17 16 15 17
Bennu Heterogeneous
N 500 50000
Error (%) 10 0 10 0
Model
POLY L inf inf inf 1
SH S 1 1 1 6
PINN III S 4 4 2 4
PINN III L 4 2 4 4
POLY S inf inf inf 5
PINN II L 5 5 6 7
PINN II S 6 6 5 8
MASCONS L 15 3 7 2
PM - 2 7 8 11
PINN I L 8 8 9 9
MASCONS S 13 10 3 10
PINN I S 9 12 10 12
SH L 7 9 12 13
GEONET L 10 11 11 14
GEONET S 11 13 13 15
ELM L 12 14 16 18
ELM S 14 15 17 19
TNN L 17 16 14 17
TNN S 16 17 15 16
Bennu Homogeneous
Table 3: Summarized rank for all models across four test asteroids
@@
While Tables 2 and 3(a) highlight the relative accuracy of the available gravity models, another important comparison point is each model’s evaluation speed. To characterize this, 1,000 randomly distributed test points are evaluated using each gravity model. The total evaluation time is measured, and an average inference time per sample is reported in Figure 7.
@@
For the large models, Figure 7 shows that the large GeodesyNet (33,601 parameters) and the large polyhedral model (30,006 parameters / approximately 10,000 facets) are the slowest gravity models reporting approximately 40 and 10 ms for each evaluation respectively. The neural network models (PINN I-III and TNN) execute an order of magnitude faster, executing between 0.4 - 0.7 ms, followed by the mascon, spherical harmonic, and ELM models at 0.3 ms, 0.09 ms, and 0.08 ms respectively. For small models, the GeodesyNet is the slowest at 3 ms, followed by the remaining machine learning gravity models averaging between 0.3 and 0.6 ms. The small polyhedral model (204 parameters or 66 faces) is faster at 0.09 ms, and the mascon, spherical harmonic, and point mass models are the fastest, executing in less than 0.01 ms per evaluation.
These results highlight that speed is among the largest trade-offs for the PINN-GM-III, particularly in the small model regime. That said, the PINN-III models are not prohibitively slow, and their computational cost remains nearly constant irrespective of model size. For example, PINN-III L is two orders of magnitude larger than PINN-III S, but its computational cost is only twice as large. In contrast, the large polyhedral models cost appears to scale linearly, taking approximately 100 times longer for a model that is 100 times bigger. This suggests that the PINN-III remains a strong option if robustness and accuracy are the primary goal, and users can increase the size and performance of these models with relatively little added overhead.
@@ 6 Conclusions
Scientific machine learning and physics informed neural networks offer a compelling set of tools to address the gravity field modeling problem. Rather than using prescriptive analytic gravity models prone to various limitations, PINNs can learn convenient representations of the gravitational potential while maintaining desirable physics properties and assurances. While past machine learning gravity models have offered early glimpses into the potential advantages of these numerical solutions, this greater class of model are often susceptible to various pitfalls that had yet to be exposed or addressed. This paper highlights these failure cases for a variety of past models, and introduces design modifications within the machine learning architecture that can overcome these challenges. Taken together, these modifications form the third generation PINN gravity model, or PINN-GM-III. This model is designed to solve the problems of feature divergence, bias towards low-altitude samples, numerical instability, and extrapolation error, while also proposing a framework for fusing analytic and numerical gravity models together for enhanced modeling accuracy. While these modifications are studied exclusively on the PINN gravity model, it should be noted that many of these modifications can likely be applied to other machine learning solutions to enhance their own robustness and generalizability.
Beyond introducing the PINN-GM-III, this manuscript also proposes new evaluation metrics to more comprehensively assess the modeling accuracy of various analytic and numerical gravity models in different orbital regimes. When trained on data from both homogeneous and heterogeneous density asteroids, the PINN-GM-III is shown to achieve competitive performance over past generations and other numerical and analytic models, while also demonstrating robustness to data sparse and noisy conditions. Future work will continue to investigate design modifications that can improve model performance, with a particular emphasis on returning to large celestial bodies and investigating ways in which the high-frequency components can be learned and represented more efficiently.
@@ Appendix A PINN-GM-III Training Details
Appendix B Hyperparameter Optimization
@@ Appendix C Comments on Past Machine Learning Performance
For GeodesyNets izzoGeodesyIrregularSmall2022 , SIRENs of 9 hidden layers with 100 nodes each are used ($\approx 100^{2}*(9-1)+8*100=80,800$). Four asteroids are studied: Bennu, Churyumov-Gerasimenko, Eros, and Itokawa. In their supplementary materials (Table S4), the relative error about Eros is reported at three characteristic altitudes. At their lowest altitude, the average error is 0.571% and their highest altitude is 0.146%. In attempts to quantify error across the entire high and low altitude regime, these values are averaged for the reported value of 0.359%. The number of training data referenced is a result of the original paper sampling 1,000 data points every 10 iterations where the models were trained for 10,000 iterations, thereby equating to 1,000,000 training data.
@@ D.5 Neural Network Models
D.6 GeodesyNet
The regression of the GeodesyNets follows the direct training method outlined in Ref. izzoGeodesyIrregularSmall2022 . The direct training method is chosen over the differential training method because the former does not assume knowledge of the asteroid shape — therefore matching the data conditions for the other machine learning gravity models.
For their regression, the GeodesyNets use 300,000 quadrature points to numerically evaluate the acceleration at a field point. In the original paper, the models are trained for 10,000 iterations, where 1,000 new training data are randomly generated in the unit volume — i.e. near the body — once every 10 iterations. In this manuscript, the data is generated a priori — either 500 or 50,000 points spanning 0-10R — and 1,000 data points are randomly sampled from this generated data once every 10 iterations to mimic the original training configuration. The architecture of the GeodesyNet matched that of the other machine learning models. The small GeodesyNet had four hidden layers and eight nodes per hidden layer, and the large GeodesyNet had eight hidden layers with 64 nodes per hidden layer. All training data were non-dimensionalized to ensure that all of the asteroid density remains within the unit cube.
The performance of the GeodesyNets in the comparative study is believed to be the result of the small dataset size and its distribution across the training volume. As discussed, the original GeodesyNet training paper achieves high accuracy by training on 1,000,000 data that exist within the unit cube or sphere; however, the data used in the comparative study experiments span a spherical volume out to 10R. This yields a volume that is approximately 1,000 times larger (volume scales cubically) with considerably less data occupying that volume. As such, the GeodesyNets do not have enough data to regress an accurate density field across the entire unit volume.
D.7 Supplemental Comparison Study Results
The following tables report the exact metric values used for the comparison study, as well as the associated rank across all models.
D.7.1 Eros Heterogeneous Density
Model
MASCONS L 50000 0 14 1.2E-2 0.13 5.4E-4 3.4E-4 8.3 1.7E-3
PINN III S 50000 0 29 0.17 1.2 2.8E-2 0.065 11 4.2E-2
PINN III L 50000 0 34 1.3E-2 0.099 1.7E-3 1.3 3.2 2.7E-3
PINN II L 50000 0 84 4.7E-2 0.080 0.070 3.2E14 2.5 2.8E-2
PINN III S 500 0 88 2.2 18 0.51 0.13 35 0.89
PINN III L 50000 10 89 3.4 3.7 2.0 0.28 13 6.3
PINN II S 50000 0 97 0.22 0.57 0.31 2.1E7 6.2 0.11
PINN III L 500 0 98 1.6 17 0.23 0.41 52 0.67
MASCONS L 500 0 101 0.82 14 2.5E-4 9.8E-8 350 1.3E-7
POLY S 110 6.2 16 3.3 0.34 23 1.9
MASCONS L 50000 10 113 1.3 9.3 0.16 0.15 610 0.37
PINN II L 500 0 115 0.69 5.1 0.22 3.2E12 25 0.18
MASCONS S 50000 0 117 3.1 4.4 2.8 2.7 26 12
POLY L 119 6.5 16 3.5 0.35 23 2.3
PINN II L 50000 10 122 1.4 2.7 2.2 4.4E14 8.8 0.68
MASCONS S 50000 10 124 3.3 4.6 3.0 3.0 26 13
MASCONS S 500 0 135 2.7 12 1.9 1.8 71 6.7
PINN III L 500 10 140 6.2 23 3.1 0.36 45 6.0
PINN II S 50000 10 140 1.5 2.9 1.8 1.1E8 9.4 6.4
PINN II S 500 0 152 1.8 7.8 1.7 5.4E6 32 3.2
PINN III S 500 10 159 8.2 51 3.3 0.36 76 5.0
PINN III S 50000 10 160 8.1 51 2.3 0.32 79 17
PINN I L 50000 0 174 10 4.3 8.4 1.6E5 8.4 12
PINN II L 500 10 179 5.7 13 4.8 5.6E11 32 5.4
PINN II S 500 10 182 6.4 13 6.5 8.6E6 33 2.7
PINN I L 50000 10 183 13 5.4 16 1.8E5 14 7.0
SH S 50000 0 202 24 370 0.45 0.22 1900 1.1
PINN I S 50000 10 202 14 9.5 31 1.6E5 31 20
PINN I S 50000 0 203 16 8.6 38 1.2E5 33 17
SH S 50000 10 211 24 320 0.53 0.29 2200 1.4
PM - 500 0 211 50 43 50 50 70 110
PM - 500 10 215 50 43 51 51 70 110
SH S 500 0 217 17 340 1.2 0.86 950 0.97
PM - 50000 10 220 53 45 53 53 70 110
PINN I L 500 10 221 29 16 48 46000 34 22
PM - 50000 0 224 53 45 53 53 70 110
SH S 500 10 227 27 580 1.6 0.74 1500 2.4
PINN I S 500 10 235 33 21 69 46000 37 31
SH L 50000 0 239 inf 1.3E99 0.45 0.22 8.8E116 1.1
PINN I S 500 0 242 41 24 75 93000 41 42
GEONET L 500 0 245 83 84 83 83 87 160
SH L 50000 10 247 inf 1.8E99 0.53 0.29 1.2E117 1.4
MASCONS S 500 10 251 34 98 29 28 460 64
GEONET L 50000 0 252 87 87 87 87 87 160
GEONET L 500 10 255 87 90 86 86 93 160
GEONET L 50000 10 260 87 87 87 87 87 160
GEONET S 500 0 273 99 99 99 99 99 180
SH L 500 10 275 inf 2.5E97 5.0 0.35 1.0E115 4.4
SH L 500 0 278 inf 9.8E96 5.0 0.35 2.1E115 4.4
PINN I L 500 0 280 52 41 100 1.1E5 47 89
GEONET S 50000 10 281 99 99 99 99 99 180
GEONET S 500 10 288 99 99 99 99 99 180
GEONET S 50000 0 292 99 99 99 99 99 180
ELM L 500 0 308 76 97 300 8.0E7 99 65
ELM L 500 10 308 78 97 310 8.4E7 99 65
ELM S 500 0 315 74 98 250 2.0E7 99 66
ELM S 500 10 315 76 98 250 2.1E7 99 66
MASCONS L 500 10 369 3.6E11 5.5E12 1.7E8 0.63 1.4E14 inf
TNN L 500 10 408 570 1200 810 56000 1100 420
TNN S 500 0 409 1300 1100 2600 1.5E5 990 450
TNN S 50000 0 410 630 1100 1900 1.7E5 1100 570
TNN L 50000 0 411 550 1100 1800 3.4E5 1600 630
TNN S 500 10 412 770 340 1200 66000 240 630
TNN S 50000 10 413 640 1100 2300 2.2E5 1200 720
TNN L 50000 10 414 610 1100 2300 2.2E5 1200 800
TNN L 500 0 415 8900 1100 6300 3.2E5 810 930
ELM L 50000 10 416 9200 1000 54000 2.6E9 520 44000
ELM L 50000 0 417 9600 1100 56000 2.6E9 550 49000
ELM S 50000 10 418 9200 1000 54000 1.8E8 520 1.9E6
ELM S 50000 0 419 9700 1100 57000 1.9E8 550 2.0E6
Table 5: Metric values for each gravity model trained on heterogeneous Eros gravity field data.